# Prompto: An open source library for asynchronous querying of LLM endpoints

### v1

@@ Abstract
Recent surge in Large Language Model (LLM) availability has opened exciting avenues for research. However, efficiently interacting with these models presents a significant hurdle since LLMs often reside on proprietary or self-hosted API endpoints, each requiring custom code for interaction. Conducting comparative studies between different models can therefore be time-consuming and necessitate significant engineering effort, hindering research efficiency and reproducibility. To address these challenges, we present prompto, an open source Python library which facilitates asynchronous querying of LLM endpoints enabling researchers to interact with multiple LLMs concurrently, while maximising efficiency and utilising individual rate limits. Our library empowers researchers and developers to interact with LLMs more effectively and enabling faster experimentation and evaluation. prompto is released with an introductory video [^1] https://youtu.be/-eZAmlV4ypk under MIT License and is available via GitHub [^2] https://github.com/alan-turing-institute/prompto.
@@
Ryan Sze-Yin Chan1, Federico Nanni1, Edwin Brown1,3, Ed Chapman1, Angus R. Williams2 Jonathan Bright2, Evelina Gabasova1 1 Research Engineering Group & 2 Public Policy Programme (The Alan Turing Institute), 3 Research Software Engineering Team (University of Sheffield) Corresponding authors: {rchan,fnanni,egabasova}@turing.ac.uk
@@ 1 Introduction
While user-friendly interfaces like ChatGPT [^3] https://chatgpt.com/, Gemini [^4] https://gemini.google.com and Claude [^5] https://claude.ai/ have made LLMs more accessible to the public, researchers typically interact with models through programmatic interfaces and APIs for greater control over parameters, automation of workflows and experiment design.
@@
Furthermore, accessing LLMs through APIs should assist reproducibility as it is possible to share code and explicit input specifications. However, La Malfa et al. (2023) notes in practice, there are several issues with the Language-Models-as-a-Service (LMaaS) paradigm (Sun et al., 2022), which refers to a licensing model in which language models are centrally hosted and typically provided on a subscription or pay-per-use basis, e.g. the OpenAI API [^6] https://openai.com/api/ and Google’s Gemini API [^7] https://ai.google.dev/gemini-api. With regards to reproducibility, since LMaaS are deployed and updated in a continuous deployment regime, models are frequently replaced by newer ones while older models can be depreciated altogether. There are further issues related to accessibility, reliability and trustworthiness (La Malfa et al., 2023). Despite these constraints, Biderman et al. (2024) suggested several best practices when evaluating LLMs such as sharing your exact prompts, parameter inputs and code and always providing model outputs.
@@
In contrast to LMaaS, there are several open LLMs, defined here as those with broadly available model weights as in Kapoor et al. (2024), e.g. Llama (Llama Team, 2024), Gemma (Gemma Team, 2024), Aya (Aryabumi et al., 2024). These "open" models are often accompanied by a model card (Wolf et al., 2020) which provides details on the instructions on executing the models locally and to provide full access to model’s internals and weights. This allows users to create their own API endpoints for LLMs on their available hardware.
@@
A significant challenge with LLMs residing on various APIs is that conducting a comparative study necessitates writing separate code to interact with each API. A further challenge is that different APIs may have different query-per-minute (QPM) rate limits. Writing separate code for each API efficiently is a time consuming engineering challenge.
@@
To address these limitations, in this system demonstration paper we introduce prompto, an open source library written in Python for asynchronous querying of multiple LLM endpoints in a consistent and highly efficient manner. prompto uses asynchronous programming to efficiently interact with endpoints by allowing users to send multiple requests to different APIs concurrently. This eliminates idle wait times and maximises efficiency, especially when dealing with different rate limits imposed by different services. In contrast, in traditional synchronous programming, a user sends a single request to a specific LLM endpoint and then waits for a response from the API before sending another request, repeating for each query. prompto currently supports a range of LMaaS endpoints (e.g. OpenAI, Gemini, Anthropic) as well as self-hosted endpoints (e.g. Ollama [^8] https://ollama.com/, Huggingface’s text-generation-interface [^9] https://github.com/huggingface/text-generation-inference library). The codebase is easily extensible to integrate new APIs and models.
@@
The library promotes experiment reproducibility by facilitating the definition of all inputs/prompts within a single JSON Lines (JSONL [^10] https://jsonlines.org/) file. The file can encompass queries for various APIs and models, enabling parallel processing for even greater efficiency gains. The library’s scalability allows it to handle large-scale experiments.
@@
In this paper, we present an overview of prompto, highlighting its modular design and flexibility. We present prompto’s functionalities, design choices, technical implementation and show its advantages in comparison with alternative approaches. We accompany its release with a series of Jupyter Notebooks as tutorials, to allow the research community to easily explore all its functionalities.
@@ 2 Related Work and Motivation
To support transparent and reproducible evaluation of large language models, there are existing evaluation frameworks such as Language Model Evaluation Harness (lm-eval [^11] https://github.com/EleutherAI/lm-evaluation-harness) (Gao et al., 2023; Biderman et al., 2024), the UK AI Safety Institute’s Inspect framework (inspect-ai [^12] https://github.com/UKGovernmentBEIS/inspect_ai), Confident AI’s DeepEval (deeepeval [^13] https://github.com/confident-ai/deepeval), LLM Comparator (llm-comparator [^14] https://github.com/pair-code/llm-comparator). These tools often have components specifically for evaluation such as automatic pipelines for model graded evaluation such as LLM-as-a-judge (Zheng et al., 2024).
@@
Rather than focusing on evaluation, our prompto library prioritises flexibility in experiment design allowing users to specify queries for various APIs and models within a single JSONL file. This not only simplifies the experimental setup for researchers, but also promotes reproducibility as all model parameters and prompts can be documented within a single file. After running an experiment, we record all model responses into an output experiment file which can be further analysed or used downstream in other pipelines.
Indeed, prompto’s applicability is intended to go beyond evaluation tasks. For instance, prompto can be leveraged to generate synthetic datasets from LLMs for model training and development (Wang et al., 2022; Xu et al., 2024a, b), or to create datasets specifically tailored for LLM distilation tasks, a technique for compressing knowledge from larger models into smaller, more efficient ones (Taori et al., 2023; Wu et al., 2023; Peng et al., 2023; Gu et al., 2023).
@@
Furthermore, our prompto library prioritises efficiency. When interacting with LLMs, a sequential and synchronous approach where requests are sent to a specific LLM endpoint and wait for a response is inefficient, especially when dealing with rate limits or the need to conduct comparisons across multiple LLMs. prompto and tackles this challenge by leveraging asynchronous programming.
@@
The ability to interact with multiple LLMs concurrently and process requests asynchronously makes our library particularly well-suited for initial exploration and rapid comparison between different LLMs. Researchers can efficiently experiment with various models and gauge their performance on specific tasks. This might be particularly useful in experiments where an "expected" or "ideal" response is not known ahead of time since evaluation framework tools are often most useful for settings where there is pre-defined model response from which the performance of an LLM can be scored against (e.g. automated metrics such as BLEU (Papineni et al., 2002) or MMLU (Hendrycks et al., 2021)). Another use case of where prompto may be preferred is in settings where the responses are intended for human evaluation.
@@ 3 Library Design
Our prompto library is an open source pip installable package [^15] https://pypi.org/project/prompto/ which facilitates processing of experiments of large language models stored as JSONL files. The library has commands to process an experiment file to query models and store results. A user is able to query multiple models in a single experiment file and prompto will query the models asynchronously and in parallel.
In this section, we focus on detailing the key components of setting up and experiment file and core commands of the library to run experiments. For a comprehensive exploration of all library features and functionalities, we kindly refer readers to the detailed documentation available within the project repository at https://alan-turing-institute.github.io/prompto/.
@@ 3.1 Setting up a prompto experiment
With prompto, an experiment file is a JSON Lines (JSONL) file which contains the prompts for an experiment along with any other parameters or metadata that is required. Each line in the JSONL file is a valid JSON [^16] https://www.json.org/ value [^17] In practice, this means each line is a collection of key/value pairs and in Python, this is realised as a dictionary. which defines a particular input to a model. We refer to a single line in the JSONL file as a prompt dictionary (or prompt_dict) with several required keys:
@@
The value of the prompt key is typically a string type that is passed to the model to generate a response, but for certain APIs or models, this could take different forms. For instance, for some API endpoints like OpenAI and Gemini, this could be a list of strings which we consider as a sequence of prompts to be sent as user messages in a chat interaction, or it could be a list of dictionaries each with "role" and "content" keys which can be used to define a history of a conversation.
There are also optional keys that can be included in a prompt dictionary such as "parameters" which define the parameters or generation configuration for a prompt such as the temperature. A "group" key could be defined to pass a user-specified grouping of the prompts which can be useful to process groups of prompts in parallel. There may also be other API or model-specific keys, for instance, with OpenAI and Gemini APIs, we allow a "multimedia" key for constructing prompts and inputs to multimodal models.
3.2 Running an experiment in prompto
  • input: contains input experiment JSONL files as described in 3.1
  • output: contains results of experiment runs
When an experiment is ran, a folder will be created in the output folder with the experiment name (the experiment file name without the ".jsonl" extension). We move the input file (originally is stored in the input folder) into the output folder and rename it to indicate it was the original input file to a particular run. The model responses are stored in a new "completed" JSONL file where each prompt dictionary will have a new "response" key storing the model response. A log file is also created to store any logs from the experiment run. Each of these files are timestamped of when the experiment started to run allowing users to run multiple times the same experiment without interfering with the same output files and to also track any changes to the input files between runs. An illustration of this data folder can be found in Figure 1.
An experiment can be ran in the terminal using the prompto_run_experiment command line interface (CLI) by specifying the path to input experiment file with the --file or -f argument. The user can also specify the path to the data folder with the --data-folder or -d argument [^18] By default, we assume the data folder is called data/ in the current working directory.:
@@ 3.2 Running an experiment in prompto
    --data-folder dataFor interacting with LLM endpoints, we often need to specify API keys or other variables. These can be set as environment variables, or they can be specified in a .env file [^19] By default, we look for a .env file in the current working directory, but a path can be specified using the --env or -e flag. as key-value pairs:
  OPENAI_API_KEY=...  GEMINI_API_KEY=...The prompto_run_experiment CLI then starts asynchronously sending requests for each prompt dictionary which includes details of the input, the API to send to and the model name to query. There are several other arguments to the command for providing flexibility of how this process occurs [^20] Full details of each of these arguments can be found in our documentation at https://alan-turing-institute.github.io/prompto/docs/pipeline/ such as the maximum queries to send per minute (--max-queries or -mq), the maximum number of retries if errors (such as rate limit or connection errors) occur when querying (--max-attempts or -ma). For example, to run an experiment where we asynchronously send 50 queries per minute (i.e. one query every 1.2 seconds) with 5 maximum retries, we can simply run the following command:
@@
    --max-attempts 5Furthermore, prompto supports sending requests to different APIs or models in parallel (--parallel or -p). However, note that our implementation of the "parallel" processing is again using asynchronous programming allowing multiple queues of prompts to be managed concurrently within a single thread of execution rather than utilising multiple CPU cores. If parallel processing of different APIs or models is requested, the user can fully customise how queues of prompts are constructed and set different rate limits for each queue [^21] Full details of parallel processing can be found at https://alan-turing-institute.github.io/prompto/docs/rate_limits/.
Alternatively, users can use all of prompto’s functionalities directly in Python rather than using the CLIs. For instance, an experiment can be ran in the following way:
@@
For a more illustrative walkthrough of running experiments, the documentation includes examples of typical workflows with associated notebooks [^22] Example usage of prompto can be found at https://alan-turing-institute.github.io/prompto/examples/.
@@ 3.3 The prompto pipeline
    --max-attempts 5This command initialises the process of continually checking the input folder for new experiments to process. Consequently, there is no need to pass in a path to an experiment file to process as with prompto_run_experiment. If a new experiment is found, it is processed and the results and logs are stored in the output folder as explained in Section 3.2. The pipeline will continue to check for new experiments until the process is stopped. In the case where there are several experiments in the input folder, the pipeline will process the experiments in the order that the files were last modified.
In Python, it is possible to initialise this process in several lines:

### v2

@@ Abstract
Recent surge in Large Language Model (LLM) availability has opened exciting avenues for research. However, efficiently interacting with these models presents a significant hurdle since LLMs often reside on proprietary or self-hosted API endpoints, each requiring custom code for interaction. Conducting comparative studies between different models can therefore be time-consuming and necessitate significant engineering effort, hindering research efficiency and reproducibility. To address these challenges, we present prompto, an open source Python library which facilitates asynchronous querying of LLM endpoints enabling researchers to interact with multiple LLMs concurrently, while maximising efficiency and utilising individual rate limits. Our library empowers researchers and developers to interact with LLMs more effectively and allowing faster experimentation, data generation and evaluation. prompto is released with an introductory video [^1] https://youtu.be/lWN9hXBOLyQ under MIT License and is available via GitHub [^2] https://github.com/alan-turing-institute/prompto.
@@
Ryan Sze-Yin Chan1, Federico Nanni1, Angus R. Williams2, Edwin Brown1,3, Liam Burke-Moore2, Ed Chapman1, Kate Onslow2, Tvesha Sippy2, Jonathan Bright2, Evelina Gabasova1 1 Research Engineering Group & 2 Public Policy Programme (The Alan Turing Institute), 3 Research Software Engineering Team (University of Sheffield) Corresponding authors: {rchan,fnanni}@turing.ac.uk
@@ 1 Introduction
While user-friendly interfaces like ChatGPT, Gemini and Claude have made LLMs more accessible to the public, researchers nowadays increasingly interact with such models through programmatic interfaces and APIs. La Malfa et al. (2023) noted several reproducibility issues with this Language-Models-as-a-Service (LMaaS) paradigm (Sun et al., 2022), where language models are centrally hosted and typically provided on a subscription or pay-per-use basis (e.g., the OpenAI API [^3] https://openai.com/api/ and Google’s Gemini API [^4] https://ai.google.dev/gemini-api). To address some of them, Biderman et al. (2024) suggested a series of best practices when evaluating LLMs, such as sharing your exact prompts, parameter inputs and code and always providing model outputs.
@@
In addition to LMaaS, there are now also several open LLMs, defined here as those with broadly available model weights as in Kapoor et al. (2024), e.g. Llama (Llama Team, 2024), Gemma (Gemma Team, 2024), Aya (Aryabumi et al., 2024). These "open" models are often accompanied by a model card (Wolf et al., 2020) which provides instructions for executing them locally and for accessing their internals and weights. This allows users to create their own API endpoints for LLMs (e.g., via Ollama [^5] https://ollama.com/) on their available hardware.
@@
A significant challenge with both LMaaS and open LLMs is that conducting a comparative study across multiple models necessitates writing separate code to interact with each API and this obstacle further hinders already complex evaluation and reproducibility efforts. An additional problem is that APIs may have different constrains (e.g., query-per-minute (QPM) rate limits), adding even more complexity to the engineering design.
@@
To address these limitations and simplify large-scale comparative studies across LLMs, we introduce prompto, an open source Python library for asynchronous querying of LLM endpoints in a consistent and highly efficient manner. prompto uses asynchronous programming to efficiently interact with endpoints by allowing users to send multiple requests to different APIs concurrently. This eliminates idle wait times and maximises efficiency, especially when dealing with different rate limits. In contrast, in traditional synchronous programming, a user sends a single request to an endpoint and waits for a response from the API before sending another request, repeating for each query. prompto supports a range of LMaaS endpoints (e.g. OpenAI, Gemini, Anthropic) as well as self-hosted endpoints for querying local models (e.g. Ollama, Hugging Face’s text-generation-inference [^6] https://github.com/huggingface/text-generation-inference for serving models hosted on Hugging Face). The codebase is easily extensible to integrate new APIs and/or locally self-hosted models. For instance, we provide an example using Quart [^7] https://github.com/pallets/quart in prompto to easily set up an endpoint for inferencing local models using transformers (Wolf et al., 2020).
@@
Inspired by Biderman et al. (2024), the library promotes experiment reproducibility by facilitating the definition of all inputs/prompts within a single JSON Lines (JSONL) or CSV file. The file can encompass queries for various APIs and models, enabling parallel processing for even greater efficiency gains. The library’s scalability allows it to handle large-scale experiments. prompto also provides built-in functionalities for automatic evaluation of the obtained responses, such as allowing the user to apply scoring functions to model outputs, and model graded evaluation or LLM-as-a-judge (Zheng et al., 2024).
@@
In this paper, we present an overview of prompto, highlighting its modular design and flexibility. We present prompto’s functionalities, design choices, technical implementation and show its advantages in comparison with alternative approaches. We accompany its release with extensive documentation and a series of Jupyter Notebooks as tutorials, to allow the research community to easily explore all its functionalities. In addition to the library’s provided examples [^8] Example usage of prompto can be found at https://alan-turing-institute.github.io/prompto/examples/, we provide an illustrative showcase of using our library and compare against a synchronous approach to query several LLM endpoints in parallel in Appendix A.
@@ 2 Related Work and Motivation
To support transparent and reproducible evaluation of large language models, a series of evaluation frameworks have been published in recent years, such as Language Model Evaluation Harness (lm-eval [^9] https://github.com/EleutherAI/lm-evaluation-harness) (Gao et al., 2023; Biderman et al., 2024), the UK AI Safety Institute’s Inspect framework (inspect-ai [^10] https://github.com/UKGovernmentBEIS/inspect_ai), Confident AI’s DeepEval (deeepeval [^11] https://github.com/confident-ai/deepeval), LLM Comparator (llm-comparator [^12] https://github.com/pair-code/llm-comparator).
@@
Nevertheless, conducting evaluation experiments on multiple large language models in order to compare their performance remains an engineering challenge, as each endpoint requires dedicated code to be written. To address this, our prompto library allows users to specify queries for various APIs and models within a single JSONL file, without having to interact directly with the details of each endpoint’s infrastructure. This prioritises flexibility in experiment design, simplifies the setup for researchers and especially promotes reproducibility as all model parameters and prompts can be documented within a single file. After running an experiment, we record all model responses into an output experiment file which can be further analysed or used downstream in other pipelines.
@@
Furthermore, our prompto library prioritises efficiency when running experiments. When interacting with LLMs, a sequential and synchronous approach where requests are sent to an endpoint and then waited upon for a response is highly inefficient, especially when dealing with different rate limits or when conducting comparisons across multiple models. prompto and tackles this challenge by leveraging asynchronous programming.
@@
The ability to interact with multiple LLMs concurrently and process requests asynchronously makes our library particularly well-suited for initial exploration and rapid comparison between different LLMs. Researchers can efficiently experiment with various models and gauge their performance on specific tasks. This might be particularly useful in experiments where an "expected" or "ideal" response is not known ahead of time since evaluation framework tools are often most useful for settings where there is pre-defined model response from which the performance of an LLM can be scored against (e.g. automated metrics such as BLEU (Papineni et al., 2002), HellaSwag (Zellers et al., 2019) or MMLU (Hendrycks et al., 2021)). prompto may also be preferred in settings where the responses are intended for human evaluation, for instance in Rao et al. (2023); Rashkin et al. (2023); Dash et al. (2023); Williams et al. (2024).
Finally, prompto’s applicability goes beyond beyond evaluation tasks, and can be leveraged to generate synthetic datasets from LLMs for model training and development (Wang et al., 2022; Xu et al., 2024a, b), or to create datasets specifically tailored for LLM distilation tasks, a technique for compressing knowledge from larger models into smaller, more efficient ones (Taori et al., 2023; Wu et al., 2023; Peng et al., 2023; Gu et al., 2023).
@@ 3 Library Design
Our prompto library is an open source Python package [^13] https://pypi.org/project/prompto/ which facilitates processing of experiments of language models stored as JSONL files. The library has commands to process an experiment file to query models and store results. A user is able to query multiple models asynchronously and in parallel in a single experiment file.
We detail key components of setting up an experiment and core commands of the library. For a comprehensive exploration of prompto’s features, we kindly refer readers to the detailed documentation at https://alan-turing-institute.github.io/prompto/.
3.1 Pipeline data folder
  • input: contains input experiment JSONL files as described in 3.2
  • output: contains results of experiment runs
When an experiment is ran, a folder will be created in the output folder with the experiment name (the experiment file name without the ".jsonl" extension). We move the input file (originally is stored in the input folder) into the output folder and rename it to indicate it was the original input file to a particular run. The model responses are stored in a new "completed" JSONL file where each prompt dictionary will have a new "response" key storing the model response. A log file is also created to store any logs from the experiment run. Each of these files are timestamped to when the experiment started, allowing users to run multiple times the same experiment without interfering with the same output files and to track any changes to the input files between runs. An illustration of this data folder can be found in Figure 1.
@@ 3.2 Setting up a prompto experiment
A prompto experiment file is a JSON Lines (JSONL) file which contains the prompts for an experiment along with any other parameters or metadata. Each line in the JSONL file is a valid JSON value [^14] In practice, this means each line is a collection of key/value pairs and in Python, this is realised as a dictionary. which defines a particular input to a model. We refer to a single line in the JSONL file as a prompt dictionary (or prompt_dict) with keys:
@@
The value of the prompt key is typically a string type that is passed to the model to generate a response, but for certain APIs or models, this could take different forms. For instance, for some API endpoints like OpenAI and Gemini, this could be a list of strings which we consider as a sequence of prompts to be sent as user messages in a chat interaction, or it could be a list of dictionaries each with "role" and "content" keys which can be used to define a history of a conversation. To foster the use of multimodal LLMs, prompto supports images and videos as inputs for several models. For instance, for the OpenAI API, it is possible to prompt with images by passing them through the "content" keys. These can be URL links to images on the web, or be stored locally in the media folder of the pipeline data folder (see Section 3.1) [^15] A full example with multimodal prompting in prompto with the OpenAI API can be found at https://alan-turing-institute.github.io/prompto/examples/openai/openai-multimodal/..
There are also optional keys that can be included in a prompt dictionary such as "parameters" which define the parameters or generation configuration for a prompt such as the temperature. Note that API services often have different names for the same generation parameter [^16] For example, to specify the maximum output tokens to generate, the OpenAI API uses max_tokens while Gemini API uses max_output_tokens.. In prompto, these parameters are not unified and the generation parameters specified in the prompt dictionary are directly passed to the API used. A "group" key could be defined to pass a user-specified grouping of the prompts which can be useful to process groups of prompts in parallel.
Note that CSV files can also be used as input where they get converted to JSONL to be processed. In this setting, the keys in the prompt dictionary discussed above are columns in the CSV file.
An experiment can be ran in the terminal using the prompto_run_experiment command line interface (CLI) by specifying the path to input experiment file with the --file or -f argument. The user can also specify the path to the data folder (Section 3.1) with the --data-folder or -d argument [^17] By default, we assume the data folder is called data/ in the current working directory.:
@@
    --data-folder dataFor interacting with LLM endpoints, we often need to specify API keys or other variables. These can be set as environment variables, or they can be specified in a .env file [^18] By default, we look for a .env file in the current working directory, but a path can be specified using --env or -e. as key-value pairs:
  OPENAI_API_KEY=...  GEMINI_API_KEY=...The prompto_run_experiment CLI then starts asynchronously sending requests for each prompt dictionary which includes details of the input, the API to send to and the model name to query. There are several other arguments to the command for providing flexibility of how this process occurs [^19] Full details of each of these arguments can be found in our documentation at https://alan-turing-institute.github.io/prompto/docs/pipeline/ such as the maximum queries to send per minute (--max-queries or -mq), the maximum number of retries if errors (such as rate limit or connection errors) occur (--max-attempts or -ma). To adhere to strict API rate limits, we use the max-queries (per minute) argument to determine how frequently to send our requests asynchronously. If max-queries is set to $1010$, we simply send our requests every 60/10=66010660/10=660 / 10 = 6 seconds. Since we use an asynchronous approach, we do not need to wait for a response before sending another request, ensuring that prompts are sent at correct intervals. To handle any errors (e.g. API failures or timeouts), we use the max-attempts argument to determine the maximum number of attempts for a prompt. In the case of an unexpected error, prompts are added to the back of the queue and are retried a maximum number of times. If the maximum is reached, we return and log the error in the response.
For example, to run an experiment where we asynchronously send 50 queries per minute (one query every 1.2 seconds) with 5 maximum retries, we can simply run the following command:
@@
    --max-attempts 5Furthermore, prompto supports sending requests to different APIs or models in parallel (--parallel or -p). Our implementation of the "parallel" processing again uses asynchronous programming allowing multiple queues of prompts to be managed concurrently within a single thread of execution rather than utilising multiple CPU cores and so is more resource efficient. If parallel processing of different APIs or models is requested, the user can fully customise how queues of prompts are constructed and set different rate limits for each queue. This is particularly useful when querying multiple API endpoints with different rate limits [^20] Full details of parallel processing can be found at https://alan-turing-institute.github.io/prompto/docs/rate_limits/.
Users can use all of prompto’s functionalities directly in Python rather than using the CLI. For instance, to run an experiment in Python:
@@
For a more illustrative walkthrough of running experiments, the documentation includes examples of typical workflows with associated notebooks\@footnotemark.
@@ 3.3 The prompto pipeline
    --max-attempts 5This command initialises the process of continually checking the input folder for new experiments to process. Consequently, there is no need to pass in a path to an experiment file to process as with prompto_run_experiment. If a new experiment is found, it is processed and the results and logs are stored in the output folder as explained in Section LABEL:subsec:experiment_run. The pipeline will continue to check for new experiments until the process is stopped. In the case where there are several experiments in the input folder, the pipeline will process the experiments in the order that the files were last modified.
In Python, it is possible to initialise this process in the following way:
@@
3.4 Automatic evaluation
A common use case for prompto is to evaluate different models, where we first need to obtain a large number of responses and then subsequently evaluate those responses. prompto provides functionality to automatically evaluate model responses.
One approach is to apply a simple scoring function to responses. A scoring function is typically lightweight such as performing string matching to some target output or some regex pattern matching. We have some built in scoring functions in the library such as a match() which determines whether the model response matches some expected response pre-defined by the user, or includes() which determines if the model responses includes a substring which is defined by the user.
In future work, we hope to expand this functionality to more advanced and computationally expensive scoring functions such toxicity classifiers (Inan et al., 2023; Hanu and Unitary team, 2020). While these can already be implemented as custom scorers in the library, scorers are applied independently to each output meaning batching is not fully utilised in the current implementation.
To automatically apply scoring functions to model outputs during an experiment run, one can simply use the --scorers argument to specify a comma-separated list of scoring functions:
  prompto_run_experiment \
    --file path/to/experiment.jsonl \
    --data-folder data \
    --scorers match,includes
Another common approach to automatic evaluation is using an LLM to judge the outputs of a model (Zheng et al., 2024). To perform an LLM-as-judge evaluation, the user must provide a prompt template which will be used to generate prompts to the Judge LLM for evaluation. In prompto, we treat this evaluation as simply as another prompto experiment where we obtain a new set of prompts using some judge evaluation template which includes a model response. An LLM-as-judge evaluation can be performed automatically after running an experiment. Full details of running LLM-as-judge evaluations and all other evaluations features available with prompto can be found at https://alan-turing-institute.github.io/prompto/docs/evaluation/.
3.5 Rephrasing prompts
It is often useful to be able to rephrase/paraphrase a given prompt, particularly in evaluation settings since performance of models can vary significantly with choice of prompt or wording (Gonen et al., 2023; Mirzadeh et al., 2024; Hughes et al., 2024). In prompto, we provide functionality to simply use another language model to rephrase a given prompt. A rephrasal experiment can be constructed by using a prompt template to prompt a model to rephrase/paraphrase. The responses from the rephrasal experiment (along with the original prompts), can be sent to another model to obtain responses for later evaluation or for any other purpose. Full details and some examples of utilising the rephrasing functionality in prompto can be found at https://alan-turing-institute.github.io/prompto/docs/rephrasals/.